Grade 11Mechanics

a helicopter of mass 2000 kg rises with a vertical acceleration of 150 m/s^2. the total mass of crew and passengers is 500 kg. then the action of rotor of the helicopter on surrounding air is?

Babzz
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
Using Newton’s second law of motion, the reaction force R’, experienced by the helicopter can be calculated as:
R’ – mg = ma
m(g + a)
= 2500 (10 + 150) = 2500 × 160
= 400000 N
The reaction force experienced by the helicopter from the surrounding air is acting upward. Hence, as per Newton’s third law of motion, the action of the rotor on the surrounding air will be 400000 N, directed downward.
 
 
Regards
Arun (askIITians forum expert)